# Memory Mosaics at scale

### v1

@@ 1 Introduction
In Machine Learning, compositional capabilities and in-context/out-of-distribution learning capabilities were continuously pursued but remain challenging. Early attempts to achieve these goals include pursuing disentanglement via various statistical “independence” (Comon, 1994; Roth et al., 2022) viewpoint, pursuing out-of-distribution/learning from the perspective of optimization on multiple environments (Finn et al., 2017; Arjovsky et al., 2019; Bengio et al., 2019). In contrast, transformer-based models demenstrate certain compositional capabilities and early in-context learning abilities. However, we still lack a clear understanding of how current transformers achieve these capabilities, and why earlier models were unable to.
@@
The contribution of this work: 1) We successfully scale up memory mosaics to llama-8B scale, using one trillion real-world training tokens. The resulting network is named as Memory Mosaics v2.111For clarity, “Memory Mosaics” refers to the version in Zhang et al. (2025), while “Memory Mosaics v2” refers to this scaled-up version. Compared to memory mosaics, memory mosaics v2 made three architectural modifications, including an adaptive bandwidth of associative memory, a gated time-variant key feature extractor, and a 3-level memory design. 2) We propose three evaluation dimensions to comprehensively assess model ability (from i.i.d. to o.o.d. scenarios). 3) Our memory mosaics v2 demonstrate superior new-task learning capabilities (with fewer examples and less priori knowledge from human designers).
This paper is organized as follows. Section 2 introduces background knowledge on associative memories. Section 3 presents the architecture of memory mosaics v2. Then section 4 describes the training process, section 5 evaluates memory mosaics v2 and transformers across three dimensions – training (persistent) knowledge storage, new knowledge storage, and in-context learning. Section 6 discusses the failure of replicating memory mosaics v2 by simply increasing the training data ($\times$8 more data) for transformers. Finally, section 7 provides discussion and further directions.
@@ 2 Background on associative memory
Associative memories have a long history in both psychology and computer science, referring to relationships between unrelated items. In this work, we follow the definition from Zhang et al. (2025), according to which an associative memory is a device that can store key-value pairs $\{(k_{1},v_{1})\dots(k_{n},v_{n})\}$ and retrieve values given a corresponding key:222both keys and values shall be vectors in $\mathbb{R}^{d}$.
@@ 3 Memory Mosaics v2
Transformers show an interesting induction head mechanism (Olsson et al., 2022), that is, predict $b$ after sequence $[\dots,a,b,\dots,a]$. This mechanism contributes to their in-context learning ability. According to Bietti et al. (2023), position encoding and asymmetric query-key extractors (e.g. $q_{T}=W_{q}x_{T}$, $k_{T}=W_{k}x_{T}$) are essential for transformers to achieve this induction head mechanism with at least two layers of attention.
@@ 3.1 Adaptive bandwidth in Gaussian kernel smoothing
where $\beta_{0}>0$, $\beta_{1}>0$, $01>\alpha>$ 0 are learnable parameters. [^5] The constraints of $\beta_{0},\beta_{1},\alpha$ are implemented as reparameterization: $\beta_{0}=e^{\theta_{0}},\beta_{1}=e^{\theta_{1}},\alpha=e^{-|\theta_{\alpha}|}$. I.e., the more key-value pairs (examples), the smaller bandwidth $1/\sqrt{\beta}$.
@@ 3.2 Gated time-variant key feature extractor
Memory mosaics (Zhang et al., 2025) employ a simple time-invariant leaky averaging to extract key features:
$k_{T}$ $=\mathrm{Norm}\big{(}{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T}}\big{)}$ $\text{with}\quad{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T}}={\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{k}_{T}}+\lambda\,{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T-1}}\quad\quad{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{k}_{T}}=W_{\varphi}\,x_{T}}\,,$ (7)
where the averaging weights $\lambda$ are fixed and independent of the semantic input $x$. As a result, semantically similar cases, such as “tom-and-jerry” and “tom- - -and- - -jerry”, may receive different key features. Inspired by recurrent-style networks (Peng et al., 2023; Gu and Dao, 2023; Beck et al., 2025), memory mosaics v2 utilize the following gated time-variant key feature extractor: [^6] It worth noting that this work focus neither on the linearization of attention nor attention efficiency. The recurrent feature extractor in Equation 7 is used to create keys, while associative memory in Equation 1 still stores all key-value pairs.
$k_{T}$ $=\mathrm{Norm}\big{(}{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T}}\big{)}\quad\text{with}\quad\begin{cases}{\,{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T}}=g_{T}{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{k}_{T}}+\lambda_{T}{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T-1}}\quad\quad{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{k}_{T}}=W_{\varphi}\,x_{T}}\\ \,g_{t}=e^{W_{g}x_{T}}\in\mathbb{R}\,,\,\lambda_{T}=e^{-|W_{\lambda}x_{T}|}\in\mathbb{R}\end{cases},$ (8)
where $W_{\varphi},W_{g},W_{\lambda}$ are learnable parameters, the averaging weights $\lambda_{T}\in\mathbb{R}$ and the the exponential gate $g_{T}\in\mathbb{R}$ semantically depend on input $x_{T}$. Check appendix figure 7 for graphical illustration.
@@
$v_{T}$ $=\mathrm{Norm}\big{(}{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{v}_{T}}\big{)}\quad\,\text{with}\quad{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{v}_{T}}=\gamma\,{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{v}_{T}}+(1-\gamma)\,{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{v}_{T+1}}\quad~{}{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{v}_{T}}=W_{\psi}\,x_{T}\,,}$ (9)
where $\gamma\in\mathbb{R}$ and $W_{\psi}$ are learnable parameters.
@@
Figure 1: Average attention scores of the last token attending previous tokens. Left: Transformer with RoPE position encoding. Right: Memory Mosaics Zhang et al. (2025). The (averaged) attention scores in transformer heavily depends on token positions (curly curves), while the attention scores in memory mosaics at far tokens (e.g. position 0 to 450) are almost invariant to positions (flat curves).
Compared with transformers (figure 1 left), the attention scores in memory mosaics exhibit a structured pattern (figure 1 right). That is, attention scores on near-tokens (positions) heavily depends on positions, while attention scores on far-tokens are almost invariant to token positions. Inspired by this experimental discovery, memory mosaics v2 replace contextual associative memory in memory mosaics with two associative memories, short-term memory and long-term memory, using distinct parameters (as figure 2 right). The short-term memory at position $t$ only store key-value pairs of near tokens, ranging from $t-h+1$ to $t-1$. In contrast, the long-term memory skips near tokens and only store key-value pairs before position $t-m$. By setting $m<h$, memory mosaics v2 create an overlap between long-term and short-term memory, resulting in a soft boundary between these two memories.
Memory mosaics v2 implements persistent memory using dense two-layers neural networks with SwiGLU activation (Shazeer, 2020) due to computational efficiency concerns. [^7] A two-layers feed-forward network and a key-value associative memory are interchangeable as shown in Sukhbaatar et al. (2019).
@@ 4 Training
We train two memory mosaics v2 of difference sizes (small/large). Memory mosaics v2 small (llama-1.5b scale) contains 24 layers, 2048 hidden dimensions, and 16 heads, trained on 200 billion tokens of a diverse datamix. Memory mosaics v2 large (llama-8b scale) increases the number of layers to 32, hidden dimensions to 4096, and the number of heads to 32, trained on 1 trillion tokens of the same datamix. Both models are trained on 4,096 context length, followed by a fine-tuning process on 32,768 context length. Other training details are provides in appendix D.
@@
Similarly, we train two baseline transformers (small/large) with the same configurations as their memory mosaics v2 counterparts. Unless otherwise specified, in this work, transformer models use llama architecture (Grattafiori et al., 2024) with multi-head attention.
@@ 5 Three evaluation dimensions
The evaluation design provides a means to assess the specific properties of a system. Memory mosaics v2 aim at the ability to learn new tasks with fewer examples and less task-specific priori knowledge. Thus, to comprehensively assess this capability, this section adopts three evaluation dimensions:
@@ 5.1 Persistent-knowledge storage and retrieval
memory mosaics v2 small 32k 35.0 60.0 58.4 32.9 73.3 62.7 58.0 11.8 46.6 29.3 3.1 34.7 30.8 59.3 27.3 9.4 49.2 1.1 38.4 38.0
@@
Table 2: Memory mosaics v2 performance on 13 common language benchmarks. Removing the “long-term memory” after training barely hurt the performance (56.6% vs 56.8%). Flops/token is estimated at context length 256 via Casson (2023) approach.
@@
Table 2 summarizes the size of parameters and computation required for transformers and memory mosaics v2. Compared to transformers, memory mosaics v2 use slightly more parameters and computations to explicitly allocate 3-level memories (persistent-memory, long-term memory, and short-term memory). It is worth paying more parameters in exchange for the 3-level memory design, because this design helps in allocating features according to their level of invariance, facilitating feature reuse in new tasks. Furthermore, removing long-term memory from memory mosaics v2 after training achieves a comparable transformer performance on the 13 persistent-knowledge benchmarks, while using fewer parameters and computations.
@@ 5.2 New-knowledge storage and retrieval
To assess this ability, we employ two “multi-unrelated-documents question-answering” tasks from the ruler benchmark Hsieh et al. (2024). These tasks involve multiple concatenated realistic articles followed by a question related to one of these articles, requiring the model to find the correct answer based on the correct article. [^8] Similarly to the process used in section 5.1 for verifying persistent-knowledge storage and retrieval tasks, appendix Table 7 compares memory mosaics v2 with and without long-term memory on these question-answering tasks, confirming the necessity of “long-term memory” for these tasks. A prompt example is:
@@
These tasks are notably more challenging than typical “needle-in-a-haystack” benchmarks (Kamradt, 2023), owing to their high information entropy.
@@
Table 3 compares memory mosaics v2 and transformers, pretrained on a 4k context length, on these question-answer tasks. Memory mosaics v2 outperforms transformers on 4k task-length by 1.4$\sim$5.6%. Similarly, Table 4 presents the same comparison, but with both models fine-tuned at a 32k context length. As task lengths increase to 32k, the “multi-unrelated-documents question-answering” tasks become more challenging. At this increased difficulty level, memory mosaics v2 significantly outperforms transformers by 12.3% to 14.8%. Moreover, memory mosaics v2 also outperforms many other public base models of similar scale (See Appendix E for the details).
@@
Table 4: Comparison of memory mosaics v2 and transformer, trained on 4k and fine-tuned on 32k context length, on ruler question-answer tasks. Memory mosaics v2 outperforms transformer by 12.3$\sim$14.8%.
@@
Many memory compression algorithms, such as rnns, xlstm (Beck et al., 2025), rwkv (Peng et al., 2023), and state-space models (Gu and Dao, 2023), fail on this task by construction because they cannot store all articles before reading the question. Similarly, local-window memory approaches, such as Alibi position encoding Press et al. (2021) and sliding-window attention Beltagy et al. (2020), also struggle for the same reason. One might argue to play around this shortage by reading the question before the multiple articles. However, this process involves task-specific priori knowledge from human designers. In the end, instead of proving the machine is intelligent, it often proves that human designers are intelligent. [^9] Please recall that a child does not prepare all questions before going to school.
@@
[…]

### v3

@@ 1 Introduction
In Machine Learning, compositional capabilities and in-context/out-of-distribution learning capabilities have been continuously pursued but remain challenging. Early attempts to achieve these goals include pursuing disentanglement via various statistical “independence” (Comon, 1994; Roth et al., 2022), pursuing out-of-distribution/learning from the perspective of optimization on multiple environments (Finn et al., 2017; Arjovsky et al., 2019; Bengio et al., 2019). In contrast, transformer-based models demonstrate certain compositional capabilities and early in-context learning abilities. However, we still lack a clear understanding of how current transformers achieve these capabilities, and why earlier models were unable to.
@@
The contribution of this work: 1) We successfully scale up memory mosaics to llama-8B scale, using one trillion real-world training tokens. The resulting network is named as Memory Mosaics v2.111For clarity, “Memory Mosaics” refers to the version in Zhang et al. (2025), while “Memory Mosaics v2” refers to this scaled-up version. [^2] https://github.com/facebookresearch/MemoryMosaics Compared to memory mosaics, memory mosaics v2 made three architectural modifications, including an adaptive bandwidth of associative memory, a gated time-variant key feature extractor, and a 3-level memory design. 2) We propose three evaluation dimensions to comprehensively assess model ability (from i.i.d. to o.o.d. scenarios). 3) Our memory mosaics v2 demonstrate superior new-task learning capabilities (with fewer examples and less priori knowledge from human designers).
This paper is organized as follows. Section 2 introduces background knowledge on associative memories. Section 3 presents the architecture of memory mosaics v2. Then section 4 describes the training process, section 5 evaluates memory mosaics v2 and transformers across three dimensions – training (persistent) knowledge storage, new knowledge storage, and in-context learning. Section 6 discusses the failure of replicating memory mosaics v2 by simply increasing the training data ($\times$8 more data) for transformers. Section 7 studies the advantage of memory mosaics v2 in fine-tuning. Finally, section 8 provides discussion and further directions.
@@ 2 Background on Associative Memory
Associative memories have a long history in both psychology and computer science, referring to relationships between unrelated items. In this work, we follow the definition from Zhang et al. (2025), according to which an associative memory is a device that can store key-value pairs $\{(k_{1},v_{1})\dots(k_{n},v_{n})\}$ and retrieve values given a corresponding key:333both keys and values shall be vectors in $\mathbb{R}^{d}$.
@@ 3 Memory Mosaics v2
Transformers show an interesting induction head mechanism (Olsson et al., 2022), that is, predict $b$ after sequence $\dots,a,b,\dots,a]$. This mechanism contributes to their in-context learning ability. According to Bietti et al. (2023), position encoding and asymmetric query-key extractors (e.g. $q_{T}=W_{q}x_{T}$, $k_{T}=W_{k}x_{T}$) are essential for transformers to achieve this induction head mechanism with at least two layers of attention.
@@ 3.1 Adaptive bandwidth in Gaussian kernel smoothing
where $\beta_{0}\geq 0,\,\beta_{1}>0,\,1>\alpha>0$ are learnable parameters (Check Appendix Table 6 for reparameterization and initialization details). I.e., the more key-value pairs (examples), the smaller bandwidth $1/\sqrt{\beta}$.
@@ 3.2 Gated time-variant key feature extractor
Memory mosaics employs a simple time-invariant leaky averaging to extract key features:
$k_{T}$ $=\mathrm{Norm}\big({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T}}\big)$ $\text{with}\quad{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T}}={\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{k}_{T}}+\lambda\,{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{k}_{T-1}}\quad\quad{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{k}_{T}}=W_{\varphi}\,x_{T}}$ (7)
The averaging weights in Equation 7 are fixed and independent of the semantic input $x$. As a result, semantically similar cases, such as “tom-and-jerry” and “tom- - -and- - -jerry”, may receive different key features. Inspired by recurrent-style networks (Peng et al., 2023; Gu and Dao, 2023; Beck et al., 2025), memory mosaics v2 utilize the following gated time-variant key feature extractor:  [^6] It worth noting that this work is neither a linearization of attention nor attention efficiency. The recurrent feature extractor in Eq. 7 is used to create keys, while associative memory in Eq. 1 still stores all key-value pairs.
[…]
where $W_{\varphi},W_{g},W_{\lambda}$ are learnable parameters, the averaging weights $\lambda_{T}\in\mathbb{R}$ and the exponential gate $g_{T}\in\mathbb{R}$ semantically depend on input $x_{T}$. See Appendix Figure 9 for graphical illustrations.
@@
$v_{T}$ $=\alpha_{\psi}\mathrm{Norm}\big({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{v}_{T}}\big)\quad\,\text{with}\quad{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\bar{v}_{T}}=\gamma\,{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{v}_{T}}+(1-\gamma)\,{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{v}_{T+1}}\quad~{\color[rgb]{0.1015625,0.5,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1015625,0.5,0.734375}\tilde{v}_{T}}=W_{\psi}\,x_{T}\,,}$ (9)
where $\gamma,\alpha_{\psi}\in\mathbb{R}$ and $W_{\psi}$ are learnable parameters.
@@
Figure 1: Average attention scores of the last token attending previous tokens. Left: Transformer with RoPE position encoding. Right: Memory Mosaics (Zhang et al., 2025). The (averaged) attention scores in transformer heavily depends on token positions (curly curves), while the attention scores in memory mosaics at far tokens (e.g. position 0 to 450) are almost invariant to positions (flat curves).
Compared with transformers (Figure 1 left), the attention scores in memory mosaics (Figure 1 right) exhibit a structured pattern. That is, attention scores on near-tokens (positions) heavily depend on positions, while attention scores on far-tokens are almost invariant to token positions. Inspired by this experimental discovery, memory mosaics v2 replace each contextual associative memory in memory mosaics with two associative memories, short-term memory and long-term memory, using distinct parameters (as in Figure 2).
Short-term memory
The short-term memory at position $t$ only stores key-value pairs of near-tokens, ranging from th+1t-h+1 to t1t-1, implementing Eq. (4) as:
$f\big({k};\>\{({k}_{t-h+1},{v}_{t-h+1})\dots({k}_{t-1},{v}_{t-1})\}\big)~=~\sum_{i=t-h+1}^{t-1}~{\footnotesize\frac{e^{\,{\beta}\,{k}^{\!\top}{k}_{i}}}{\sum_{j=t-h+1}^{t-1}e^{\,{\beta}\,{k}^{\!\top}{k}_{j}}}}~{v}_{i}~.$ (10)
Long-term memory
[…]
Persistent Memory
Memory mosaics v2 implements persistent memory using dense two-layer neural networks with SwiGLU activation (Shazeer, 2020) due to computational efficiency concerns. [^8] A two-layers feed-forward network and a key-value associative memory are interchangeable as shown in Sukhbaatar et al. (2019).
@@ 4 Training
We train two memory mosaics v2 of difference sizes (small/large). Memory mosaics v2 small (llama-1.5b scale) contains 24 layers, 2048 hidden dimensions, and 16 heads, trained on 200 billion tokens of a diverse datamix. Memory mosaics v2 large (llama-8b scale) increases the number of layers to 32, hidden dimensions to 4096, and the number of heads to 32, trained on 1 trillion tokens of the same datamix. Both models are trained on 4,096 context length, followed by a fine-tuning process on 32,768 context length. Other training details are provided in Appendix C.
Stochastic long-term memory size
During training, memory mosaics v2 samples the long-term memory delay step $m$ from $[64,256]$, sets the short-term memory window size $h=256$. At inference, $m$ is set to 64. This stochastic long-term memory training setup encourages the allocation of position-invariant signals to long-term memory and position-dependent signals to short-term memory (as shown in Figure 1). The experimental results in Appendix G Table 12 show that this training setup enhances context-length extrapolation ability by more than 15%.
@@
We train two baseline transformers (small/large) with the same configurations as their memory mosaics v2 counterparts. Unless otherwise specified, in this work, transformer models use llama architecture (Grattafiori et al., 2024) with multi-head attention.
@@ 5 Three evaluation dimensions
The evaluation design provides a means to assess the specific properties of a system. Memory mosaics v2 aims at the ability to learn new tasks with fewer examples and less task-specific priori knowledge (Zhang, 2025). Thus, to fully assess this capability, this section adopts three evaluation dimensions.
@@ 5.1 Persistent-knowledge storage and retrieval
memory mosaics v2 small 32k 35.0 60.0 58.4 32.9 73.3 62.7 58.0 11.8 46.6 29.3 3.1 34.7 30.8 59.3 27.3 1.1 9.4 49.2 38.4 38.0
@@
Table 2: Memory mosaics v2 performance on 13 common language benchmarks. Removing the “long-term memory” after training barely hurt the performance (56.6% vs 56.8%). Flops/token is estimated at context length 256 via tha approach of Casson (2023).
@@
Table 2 summarizes the size of parameters and computation required for transformers and memory mosaics v2. Interestingly, removing long-term memory from memory mosaics v2 after training achieves a comparable transformer performance on the 13 persistent-knowledge benchmarks, while using fewer parameters and computations.
@@ 5.2 New-knowledge storage and retrieval
To assess this ability, we employ two “multi-unrelated-documents question-answering” tasks from the ruler benchmark (Hsieh et al., 2024). These tasks involve multiple concatenated realistic articles followed by a question related to one of these articles, requiring the model to find the correct answer based on the correct article. [^9] Similarly to the process used in section 5.1 for verifying persistent-knowledge storage and retrieval tasks, appendix Table 10 compares memory mosaics v2 with and without long-term memory on these question-answering tasks, confirming the necessity of “long-term memory” for these tasks. A prompt example is:
@@
These tasks are notably more challenging than typical ’needle-in-a-haystack’ benchmarks (Kamradt, 2023), owing to their high information entropy. The typical ’needle-in-a-haystack’ task is too easy, resulting in many models achieving near-perfect performance. See Table 13 in Appendix for details.
@@
Table 3: Comparison of memory mosaics v2 and transformer, trained on 4k context length, on ruler question-answer tasks. Memory mosaics v2 not only outperforms transformer on 4k task-length, but also successfully extrapolate the context length $\times 4\sim\times 8$ times without any fine-tuning.
@@
[…]
@@
Many memory compression algorithms, such as rnns, xlstm (Beck et al., 2025), rwkv (Peng et al., 2023), and state-space models (Gu and Dao, 2023), fail on this task by construction because they cannot store all articles before reading the question. Similarly, local-window memory approaches, such as Alibi position encoding Press et al. (2021) and sliding-window attention Beltagy et al. (2020), also struggle for the same reason. [^10] One might argue to play around this shortage by reading the question before the multiple articles. However, this process involves task-specific priori knowledge from human designers. In the end, instead of proving the machine is intelligent, it often proves that human designers are intelligent. Please recall that a child does not prepare all questions before going to school. This incompetent of memory compression algorithms has also been experimentally demonstrated by Hsieh et al. (2024) and Li et al. (2024). Also, see Appendix G for these experimental evidences.
@@
Context length extrapolation (without fine-tuning) not only is computationally appealing, but also reveals the model’s consistency in handling context. Unfortunately, transformers (with rope position encoding) struggle to extrapolate context length, as shown in Table 3.111111The comparison ignores many memory compression and local window approaches (Press et al., 2021; Beltagy et al., 2020), because they fail on this evaluation by construction. In contrast, memory mosaics v2, trained on 4k context length, not only outperform transformers on 4k length, but also perform well after extrapolating context length $\times 4\sim\times 8$ times without any fine-tuning or adaptation.